# Causality-Aware Temporal Projection for Video Understanding in Video-LLMs

### v1

@@ I Introduction
Video understanding is inherently a temporally structured reasoning problem. Unlike static images, the semantic meaning of a video emerges from ordered event transitions and their causal dependencies over time. With the rapid progress of Large Language Models (LLMs) [achiam2023gpt4, touvron2023llama2], recent research has shifted from coarse activity recognition toward more complex video question answering (VQA) tasks that require explaining why and how events unfold [liu2023llava, li2023blip2]. To enable efficient multimodal reasoning, many parameter-efficient Video-LLMs [maaz2024videochatgpt, lin2023videollava, zhang2023videollama] adopt a two-stage pipeline: a frozen visual encoder extracts frame-level features, which are then mapped into the LLM token space via a lightweight projection module. This design allows effective video-language alignment without modifying the heavy LLM backbone.
Despite their effectiveness, existing parameter-efficient Video-LLMs often lack explicit architectural mechanisms to model temporal ordering in a structured manner. To manage temporal dependencies under strict computational constraints, most approaches rely on heuristic pooling strategies or unconstrained bidirectional transformers to aggregate inter-frame information. While such designs are flexible and expressive, they may introduce temporal leakage, where information from later frames influences representations of earlier events, thereby weakening ordered temporal reasoning. As illustrated in Fig. 1(a), this effect is particularly undesirable in benchmarks like NExT-QA [xiao2021next], where reasoning about event order and causality is central to the task.
@@
To ensure practical applicability, we adopt a 4-bit QLoRA [dettmers2023qlora] training paradigm with LoRA adapters [hu2021lora]. By freezing both the visual encoder and the LLM backbone and optimizing only lightweight modules, V-CORE enables end-to-end training on a single consumer GPU. Experimental results demonstrate the effectiveness of explicit temporal ordering. On the challenging NExT-QA benchmark [xiao2021next], V-CORE achieves 61.2% accuracy, with the most notable improvements concentrated in the Temporal and Causal sub-categories, while maintaining competitive performance on descriptive questions.
@@ II-A Video-Language Models
Video-Language Models (Video-LLMs) extend the instruction-following capabilities of LLMs [achiam2023gpt4, touvron2023llama2] to the video domain by coupling a frozen vision encoder with a lightweight projection module. Early systems such as Video-ChatGPT [maaz2024videochatgpt] and Video-LLaMA [zhang2023videollama] established this paradigm using linear or shallow projectors to map visual features into the LLM token space. Subsequent works, including Video-LLaVA [lin2023videollava], further improved multimodal alignment through instruction tuning and better visual-language pretraining. More recent models such as LLaVA-NeXT [zhang2024llavanext] and Qwen2-VL [wang2024qwen2vl] scale video understanding by increasing visual granularity through dynamic resolution and advanced token compression strategies. While these approaches achieve strong performance, temporal modeling is primarily handled implicitly within the LLM’s self-attention layers. The projection stage itself is generally agnostic to temporal directionality, treating video frames as an unordered or symmetrically related set. Fine-grained multimodal understanding of complex visual content has also been studied in prior work [deng2024covis, wang2022internvideo], complementing the video temporal reasoning focus. Normally, these task-incremental settings can reveal forgetting behaviors in multimodal models, which has been benchmarked in recent work [li2025mcl]. In contrast, our work focuses on parameter-efficient video QA and introduces explicit architectural constraints at the projection stage to preserve temporal ordering, rather than relying on implicit temporal reasoning to emerge within the LLM.
@@ II-B Spatio-Temporal Modeling and Token Compression
Modeling temporal structure is central to video understanding, as it captures motion, event ordering, and causal dependencies. Vision Transformers such as TimeSformer [bertasius2021timesformer] and ViViT [arnab2021vivit] explicitly model spatio-temporal interactions, but their computational cost limits their direct adoption as connectors for large language models. Structured attention and relation modeling [kang2025lpdetr, yu2024spatial] have been explored in vision transformers to improve efficiency and representation quality. To maintain efficiency, many Video-LLMs compress frame-level features into a compact set of tokens before projection. Recent methods such as ST-LLM [liu2024stllm] and LLaMA-VID [li2024llamavid] reduce redundancy through token pruning or aggregation, but typically rely on global or bidirectional attention mechanisms within the temporal projector. As highlighted by recent benchmarks such as Video-MME [fu2024videomme], the absence of explicit directional constraints can hinder long-range temporal reasoning, particularly in tasks requiring event ordering. Unlike prior approaches that primarily focus on reducing token count, our method jointly addresses what information to preserve and how it is temporally structured. Specifically, Learnable Spatial Aggregation selects salient spatial tokens per frame, while the Causality-Aware Temporal Projector enforces a unidirectional temporal flow consistent with the arrow of time, culminating in a terminal causal sink for global aggregation.
@@ II-C Parameter-Efficient Fine-Tuning (PEFT)
Parameter-Efficient Fine-Tuning (PEFT) has become the de facto standard for adapting large backbones. LoRA [hu2021lora] and QLoRA [dettmers2023qlora] enable fine-tuning 7B+ models on consumer-grade hardware by injecting low-rank trainable matrices. While early Video-LLMs primarily trained the multimodal projector, recent methods have increasingly applied LoRA to the LLM layers to enable deeper modality alignment. Our method builds upon the QLoRA paradigm but fundamentally revisits the architecture of the temporal connector. Instead of a frame-wise linear layer that treats frames as independent entities, we employ a causality-aware temporal transformer featuring a terminal causal sink to explicitly enforce unidirectional temporal information flow. This design mitigates temporal leakage that can arise from unconstrained bidirectional attention in temporally ordered video representations. This architectural refinement ensures that the visual context provided to the LLM is both temporally ordered and causally consistent.
@@ III-F Parameter-Efficient Adaptation
We utilize QLoRA [dettmers2023qlora] to insert low-rank adapters into the attention layers of the frozen Vicuna-7B. This enables the LLM to adapt to the structured temporal tokens produced by CATP, rather than learning video representations from scratch. The model is optimized using the auto-regressive loss on general video-instruction datasets, enabling robust zero-shot generalization. This design ensures that performance gains primarily stem from the proposed spatio-temporal modeling modules rather than increased model capacity.
@@ IV-A Datasets and Metrics
We evaluate our V-CORE framework on four VideoQA benchmarks (Table I). MSVD-QA [ye2017video] and MSRVTT-QA [xu2016msr] focus on short-range visual grounding. TGIF-QA [jang2017tgif] demands sensitivity to motion patterns (Action and Transition subsets). Our primary benchmark, NExT-QA [xiao2021next], requires complex reasoning to explain event dependencies in long-form videos. Following Video-ChatGPT [maaz2024videochatgpt], we report Accuracy (Acc) and a semantic score (0–5 scale) for open-ended questions. Since video-language models may generate responses with varying lengths and expressions, traditional word matching metrics are insufficient. We therefore adopt an LLM-assisted evaluation protocol, where GPT-3.5-turbo judges the correctness and semantic relevance of predicted answers given the question and ground-truth response.
@@ IV-B Implementation Details
To ensure parameter efficiency, a 4-bit NormalFloat (NF4) QLoRA [dettmers2023qlora] is applied only to the query ($W_{q}$) and value ($W_{v}$) projection layers of the LLM. We set the LoRA rank to $r=64$ with a scaling factor $\alpha=128$. This lightweight adaptation strategy ensures that performance gains primarily stem from the proposed spatio-temporal modeling modules.
@@
Video-LLaMA [zhang2023videollama] LLaMA-7B 150K 51.6 2.5 29.6 1.8 - - - -
VISTA-LLaMA [ma2024vista] LLaMA-7B 100K 65.3 3.6 60.5 3.3 - - 60.7 3.4
Video-ChatGPT [maaz2024videochatgpt] Vicuna-7B 100K 64.9 3.3 49.3 2.8 51.4 3.0 54.6 3.2
@@
VideoChat [li2023videochat] Vicuna-7B 10M 56.3 2.8 45.0 2.5 34.4 2.3 56.2 3.2
Video-LLaVA [lin2023videollava] Vicuna-7B 700K 70.4 3.9 58.3 3.5 70.0 4.0 - -
VideoChat2 [li2024mvbench] Vicuna-7B 2M 70.0 3.9 54.1 3.3 60.4 - 61.7 -

### v2

@@ I Introduction
Video understanding is inherently a temporally structured reasoning problem. Unlike static images, the semantic meaning of a video emerges from ordered event transitions and their causal dependencies over time. With the rapid progress of Large Language Models (LLMs) [1, 21], recent research has shifted from coarse activity recognition toward more complex video question answering (VQA) tasks that require explaining why and how events unfold [17, 11]. To enable efficient multimodal reasoning, many parameter-efficient Video-LLMs [20, 16, 28] adopt a two-stage pipeline: a frozen visual encoder extracts frame-level features, which are then mapped into the LLM token space via a lightweight projection module. This design allows effective video-language alignment without modifying the heavy LLM backbone.
Despite their effectiveness, existing parameter-efficient Video-LLMs often lack explicit architectural mechanisms to model temporal ordering in a structured manner. To manage temporal dependencies under strict computational constraints, most approaches rely on heuristic pooling strategies or unconstrained bidirectional transformers to aggregate inter-frame information. While such designs are flexible and expressive, they may introduce temporal leakage, where information from later frames influences representations of earlier events, thereby weakening ordered temporal reasoning. As illustrated in Fig. 1(a), this effect is particularly undesirable in benchmarks like NExT-QA [24], where reasoning about event order and causality is central to the task.
@@
To ensure practical applicability, we adopt a 4-bit QLoRA [6] training paradigm with LoRA adapters [8]. By freezing both the visual encoder and the LLM backbone and optimizing only lightweight modules, V-CORE enables end-to-end training on a single consumer GPU. Experimental results demonstrate the effectiveness of explicit temporal ordering. On the challenging NExT-QA benchmark [24], V-CORE achieves 61.2% accuracy, with the most notable improvements concentrated in the Temporal and Causal sub-categories, while maintaining competitive performance on descriptive questions.
@@ II-A Video-Language Models
Video-Language Models (Video-LLMs) extend the instruction-following capabilities of LLMs [1, 21] to the video domain by coupling a frozen vision encoder with a lightweight projection module. Early systems such as Video-ChatGPT [20] and Video-LLaMA [28] established this paradigm using linear or shallow projectors to map visual features into the LLM token space. Subsequent works, including Video-LLaVA [16], further improved multimodal alignment through instruction tuning and better visual-language pretraining. More recent models such as LLaVA-NeXT [29] and Qwen2-VL [22] scale video understanding by increasing visual granularity through dynamic resolution and advanced token compression strategies. While these approaches achieve strong performance, temporal modeling is primarily handled implicitly within the LLM’s self-attention layers. As a result, the projection stage is often agnostic to temporal directionality, treating frame-level features as unordered or symmetrically related tokens.
Complementary to video-centric modeling, fine-grained multimodal understanding of complex visual content has been studied in prior work [5, 23], highlighting the importance of structured visual representations for downstream reasoning. Beyond accuracy, recent studies have also examined reliability aspects of multimodal models, including robustness under visual adversarial perturbations [4] and stability under task-incremental learning settings where forgetting can occur [15]. In contrast, our work focuses on parameter-efficient video QA and introduces explicit architectural constraints at the projection stage to preserve temporal ordering, rather than relying on implicit temporal reasoning to emerge within the LLM.
@@ II-B Spatio-Temporal Modeling and Token Compression
Modeling temporal structure is central to video understanding, as it captures motion, event ordering, and causal dependencies. Vision Transformers such as TimeSformer [3] and ViViT [2] explicitly model spatio-temporal interactions, but their computational cost limits their direct adoption as connectors for large language models. Structured attention and relation modeling [10, 27] have been explored in vision transformers to improve efficiency and representation quality. To maintain efficiency, many Video-LLMs compress frame-level features into a compact set of tokens before projection. Recent methods such as ST-LLM [18] and LLaMA-VID [14] reduce redundancy through token pruning or aggregation, but typically rely on global or bidirectional attention mechanisms within the temporal projector. As highlighted by recent benchmarks such as Video-MME [7], the absence of explicit directional constraints can hinder long-range temporal reasoning, particularly in tasks requiring event ordering. Unlike prior approaches that primarily focus on reducing token count, our method jointly addresses what information to preserve and how it is temporally structured. Specifically, Learnable Spatial Aggregation selects salient spatial tokens per frame, while the Causality-Aware Temporal Projector enforces a unidirectional temporal flow consistent with the arrow of time, culminating in a terminal causal sink for global aggregation.
@@ II-C Parameter-Efficient Fine-Tuning (PEFT)
Parameter-Efficient Fine-Tuning (PEFT) has become the de facto standard for adapting large backbones. LoRA [8] and QLoRA [6] enable fine-tuning 7B+ models on consumer-grade hardware by injecting low-rank trainable matrices. While early Video-LLMs primarily trained the multimodal projector, recent methods have increasingly applied LoRA to the LLM layers to enable deeper modality alignment. Our method builds upon the QLoRA paradigm but fundamentally revisits the architecture of the temporal connector. Instead of a frame-wise linear layer that treats frames as independent entities, we employ a causality-aware temporal transformer featuring a terminal causal sink to explicitly enforce unidirectional temporal information flow. This design mitigates temporal leakage that can arise from unconstrained bidirectional attention in temporally ordered video representations. This architectural refinement ensures that the visual context provided to the LLM is both temporally ordered and causally consistent.
@@ III-F Parameter-Efficient Adaptation
We utilize QLoRA [6] to insert low-rank adapters into the attention layers of the frozen Vicuna-7B. This enables the LLM to adapt to the structured temporal tokens produced by CATP, rather than learning video representations from scratch. The model is optimized using the auto-regressive loss on general video-instruction datasets, enabling robust zero-shot generalization. This design ensures that performance gains primarily stem from the proposed spatio-temporal modeling modules rather than increased model capacity.
@@ IV-A Datasets and Metrics
We evaluate our V-CORE framework on four VideoQA benchmarks (Table I). MSVD-QA [26] and MSRVTT-QA [25] focus on short-range visual grounding. TGIF-QA [9] demands sensitivity to motion patterns (Action and Transition subsets). Our primary benchmark, NExT-QA [24], requires complex reasoning to explain event dependencies in long-form videos. Following Video-ChatGPT [20], we report Accuracy (Acc) and a semantic score (0–5 scale) for open-ended questions. Since video-language models may generate responses with varying lengths and expressions, traditional word matching metrics are insufficient. We therefore adopt an LLM-assisted evaluation protocol, where GPT-3.5-turbo judges the correctness and semantic relevance of predicted answers given the question and ground-truth response.
@@ IV-B Implementation Details
To ensure parameter efficiency, a 4-bit NormalFloat (NF4) QLoRA [6] is applied only to the query ($W_{q}$) and value ($W_{v}$) projection layers of the LLM. We set the LoRA rank to $r=64$ with a scaling factor $\alpha=128$. This lightweight adaptation strategy ensures that performance gains primarily stem from the proposed spatio-temporal modeling modules.
@@
Video-LLaMA [28] LLaMA-7B 150K 51.6 2.5 29.6 1.8 - - - -
VISTA-LLaMA [19] LLaMA-7B 100K 65.3 3.6 60.5 3.3 - - 60.7 3.4
Video-ChatGPT [20] Vicuna-7B 100K 64.9 3.3 49.3 2.8 51.4 3.0 54.6 3.2
@@
VideoChat [12] Vicuna-7B 10M 56.3 2.8 45.0 2.5 34.4 2.3 56.2 3.2
Video-LLaVA [16] Vicuna-7B 700K 70.4 3.9 58.3 3.5 70.0 4.0 - -
VideoChat2 [13] Vicuna-7B 2M 70.0 3.9 54.1 3.3 60.4 - 61.7 -